### v1

# The action of an inverse semigroup on its Stone-Čech compactification

Abstract.
We initiate the study of the Stone-Čech transformation groupoid $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ of an inverse semigroup $\mathcal{S}$. This groupoid occurs as a model for the inverse semigroup Roe algebra $\mathcal{R}_{\mathcal{S}}$ of Lledo and Martinez; in particular, the action of $\mathcal{S}$ on $\beta\mathcal{S}$ is amenable in the sense of Exel and Starling if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact. We prove that the properties of being Hausdorff, principal, and effective are all equivalent for $\mathcal{G}$, and give an algebraic condition on $\mathcal{S}$ equivalent to the Hausdorffness of $\mathcal{G}$. Finally, we show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$.
1. Introduction
footnotetext: 2020 Mathematics Subject Classification Primary: 20M18, 18B40, 46L05; Secondary: 43A65.
Keywords: étale groupoids, inverse semigroups, Stone-Čech compactification.
Part of this work was completed as part of the Master’s thesis of the first author. The second author is funded by NSERC grant RGPIN-2021-03834.
A semigroup $\mathcal{S}$ is called inverse if, for every $s\in\mathcal{S}$, there is a unique $s^{*}\in\mathcal{S}$ with $ss^{*}s=s$ and $s^{*}ss^{*}=s^{*}$. The dynamics and geometry of inverse semigroups have been productively studied in a way that generalizes the discrete group setting. In [lledomartinez], Lledo and Martinez described the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ of a countable inverse semigroup $\mathcal{S}$, and showed that this C-algebra was naturally isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
In this note, we describe an étale groupoid model for $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. Like the group case, this étale groupoid arises from a left-translation action of $\mathcal{S}$ on its Stone-Čech compactification (Theorem 3.1). It follows immediately that Exel’s reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable as an étale groupoid. Said differently, the following generalization of the discrete group case ([ozawa], Theorem 3) holds:
Theorem 1.
Exel’s inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable in the sense of Exel and Starling [starling].
It is well-known that the action of a discrete group $\Gamma$ on $\beta\Gamma$ is free (see for example [NR14, Example 2.14]. In the inverse semigroup setting, one may therefore anticipate that $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ is always principal. Using a finite cover criterion of Exel and Pardo [EP16], we obtain the following generalization of this result.
Theorem 2.
Let $\mathcal{S}$ be an inverse semigroup. Then the following statements are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff,
  2. (2)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal,
  3. (3)
    every clopen action of $\mathcal{S}$ on a locally compact Hausdorff space has a Hausdorff transformation groupoid.
This generalizes the group case because $\mathcal{S}\ltimes\beta\mathcal{S}$ is always Hausdorff when $\mathcal{S}$ is a group. Using Theorem 2, we give a criterion on $\mathcal{S}$ (Theorem 3.8) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.9). Finally, since there are many examples of inverse semigroups with non-Hausdorff tight groupoids, condition (3) of Theorem 2 provides many examples of inverse semigroups for which $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff nor principal.
2. Preliminaries
In what follows, $\mathcal{S}$ denotes a countable inverse semigroup. An element $e\in\mathcal{S}$ is called idempotent if $e^{2}=e$. For any $\in\mathcal{S}, ss$$^{*}s$ and $ss^{*}$ are idempotent. Denote by $E_{\mathcal{S}}$ the set of idempotent elements of $\mathcal{S}$. It is a pleasant fact that $E_{\mathcal{S}}$ is a commutative subsemigroup of $\mathcal{S}$, and that $\mathcal{S}$ is a group if and only if $E_{\mathcal{S}}$ is a singleton. If $X$ is a set, then the collection of all partial bijections between subsets of $X$, denoted by $\mathcal{I}(X)$, is an inverse semigroup under the operation of partial composition called the symmetric inverse monoid on $X$.
Inverse semigroups carry a natural partial order: $s\leqslant t$ if and only if $ts^{*}s=s$. Restricted to $E_{\mathcal{S}}$, this order is given by $e\leqslant f$ if and only if $ef=e$.
For any poset $(P,\prec)$ and $A\subseteq P$ we will use the notation
$A^{\prec}:=\{b\in P:a\prec b\text{ for some }a\in A\}.$
In this paper, we only consider the partial orders $\subseteq$ on the power set of $\mathcal{S}$, and $\leqslant$ or $\geqslant$ on $\mathcal{S}$.
We now introduce actions of inverse semigroups on topological spaces and C-algebras. Although this note is only interested in two such actions (one on $\beta\mathcal{S}$, and one on $\ell^{\infty}(\mathcal{S})$), we feel it will be beneficial to introduce the theory in somewhat excessive generality. Let $\mathcal{A}$ be a C-algebra. An action of an inverse semigroup $\mathcal{S}$ on $\mathcal{A}$ is a semigroup homomorphism $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{A})$ such that, for each $s\in\mathcal{S}$,
  1. (1)
    The domain and codomain of $\alpha_{s}$ is a closed, two-sided ideal in $\mathcal{A}$;
  2. (2)
    The partial bijection $\alpha_{s}$ is a *-homomorphism.
Given an idempotent $e\in E_{\mathcal{S}}$, we set $I_{e}=\text{dom}(\alpha_{e})$. We observe that $\alpha_{s}$ is a *-isomorphism of $I_{s^{*}s}$ onto $I_{ss^{*}}$. The next example, central in [lledomartinez], is also the main C-action we will consider in this note.
Example 2.1.
Let $\mathcal{A}=\ell^{\infty}(\mathcal{S})$. For each $e\in E_{\mathcal{S}}$, define $I_{e}=\{f\in\mathcal{A}:\text{supp}(f)\subseteq e\mathcal{S}\}$. For each $s\in\mathcal{S}$, define $\alpha_{s}:I_{s^{*}s}\to I_{ss^{*}}$ by $\alpha_{s}f](t)=f(s^{*}t)$. The mapping $s\in\mathcal{S}\mapsto\alpha_{s}\in\mathcal{I}(\mathcal{S})$ is an action of $\mathcal{S}$ on $\mathcal{A}$.
The C-algebra $\ell^{\infty}(\mathcal{S})$ is *-isomorphic to $C(\beta\mathcal{S})$, where $\beta\mathcal{S}$ is the Stone-Čech compactification of $\mathcal{S}$. At the level of compact Hausdorff spaces, inverse semigroup actions are described as follows (see [EXEL] for details): an action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
  1. (1)
    $\alpha_{s}$ is continuous, for all $s\in\mathcal{S}$;
  2. (2)
    $\text{dom}(\alpha_{s})$ is open, for all $s\in\mathcal{S}$.
[…]
Example 2.2 ([gondek24], Section 3.2).
Let us define an inverse semigroup action of $\mathcal{S}$ on the Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding. For each $s\in\mathcal{S}$, define a bijection $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ by $\lambda_{s}t=st$ ($t\in s^{*}\mathcal{S}$). By the universal property of $\beta\mathcal{S}$, $\lambda_{s}$ extends to a homeomorphism $\lambda_{s}:\overline{s^{*}\mathcal{S}}\to\overline{s\mathcal{S}}$, where $\overline{s^{*}\mathcal{S}}:=\overline{\iota(s^{*}\mathcal{S})}^{\beta}\subseteq\beta\mathcal{S}$. We observe that the domain and codomain of $\lambda_{s}$ are always a compact, open subsets of $\beta\mathcal{S}$. It is routine to check that the mapping $\mathcal{S}\ni s\mapsto\lambda_{s}\in\mathcal{I}(\beta\mathcal{S})$ is a semigroup homomorphism, and therefore an action, which we shall henceforth refer to by $\mathcal{S}\curvearrowright\beta\mathcal{S}$.
The contravariant action $\hat{\lambda}$ of $\mathcal{S}$ on $C(\beta\mathcal{S})$ is the action in Example 2.1, after applying the identification $j:C(\beta\mathcal{S})\ni f\mapsto f\circ\iota\in\ell^{\infty}(\mathcal{S})$. Indeed, given $f\in I_{s^{*}s}\subseteq\ell^{\infty}(\mathcal{S})$, we have $j(f)](t)=f(\iota(t))=0\Leftrightarrow\iota(t)\not\in\overline{s^{*}\mathcal{S}}\Leftrightarrow t\not\in s^{*}\mathcal{S}$, so that $j(\text{dom}(\hat{\lambda}_{s}))=I_{s^{*}s}$, and
$[\alpha_{s}(j(f))](t)=[j(f)](s^{*}t)=f(\iota(s^{*}t))=f(\lambda_{s^{*}}\iota(t))=[\hat{\lambda}_{s}f](t),$
for all $f\in I_{s^{*}s}$.
Given a C-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ and an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{A}$, there is a standard construction for a reduced crossed product C-algebra $\mathcal{A}\rtimes_{r}\mathcal{S}\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$, which does not depend on the inclusion $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ (for details, see [EXEL]). We will not work directly with this C-algebra, so we only sketch the construction. For each $s\in\mathcal{S}$, we define the maps $V_{s}:\mathcal{S}\to\mathcal{B}(\ell^{2}(\mathcal{S}))$ and $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}\otimes\mathcal{\ell}^{2}(\mathcal{S}))$ by
$V_{s}\delta_{t}=\begin{cases}\delta_{st}&t\in s^{*}\mathcal{S}\\ 0&\text{otherwise}\end{cases}\qquad$ $\text{and}\qquad\pi(a)(\xi\otimes\delta_{t})=\begin{cases}\alpha_{t^{*}}(a)\xi\otimes\delta_{t}&t\in I_{s^{*}s}\\ 0&\text{otherwise}.\end{cases}$
The reduced crossed product is the C-subalgebra of $\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$ defined by
$\mathcal{A}\rtimes_{r}\mathcal{S}:=\textup{C}^{*}(\{1\otimes V_{s}:s\in\mathcal{S}\}\cup\pi(A))\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S})).$
We are interested in describing an étale groupoid $\mathcal{G}$ with the property that $\text{C}^{*}_{r}(\mathcal{G})=\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In preparation for this, let us establish conventions for étale groupoids in this note. Given a groupoid $\mathcal{G}$, we will denote by $\mathcal{G}^{(2)}\subset\mathcal{G}\times\mathcal{G}$ the composable pairs, and $\mathcal{G}^{(0)}=\{\gamma^{-1}\gamma:\gamma\in\mathcal{G}\}$ the unit space. The source and range maps $\mathbf{d}:\mathcal{G}\to\mathcal{G}^{(0)}$ and $\mathbf{r}:\mathcal{G}\to\mathcal{G}^{(0)}$ (respectively) are defined by $\mathbf{d}(\gamma)=\gamma^{-1}\gamma$ and $\mathbf{r}(\gamma)=\gamma\gamma^{-1}$, for $\gamma\in\mathcal{G}$. Given $u\in\mathcal{G}^{(0)}$, we will write $\mathcal{G}_{u}=\mathbf{d}^{-1}(u)$, $\mathcal{G}^{u}=\mathbf{r}^{-1}(u)$, and $\mathcal{G}^{u}_{u}=\mathcal{G}_{u}\cap\mathcal{G}^{u}$. The isotropy $\textup{Iso}(\mathcal{G})$ is defined by
$\textup{Iso}(\mathcal{G})=\{\gamma\in\mathcal{G}:\textbf{r}(\gamma)=\textbf{d}(\gamma)\}.$
We recall that the groupoid $\mathcal{G}$ is called principal if $\textup{Iso}(\mathcal{G})=\mathcal{G}^{(0)}$.
A topological groupoid is a groupoid $\mathcal{G}$ equipped with a topology under which the multiplication $\mathcal{G}^{(2)}\ni(\alpha,\beta)\mapsto\alpha\beta$ and inverse map $\mathcal{G}\ni\gamma\mapsto\gamma^{-1}$ are continuous. A topological groupoid is called étale if it is locally compact, $\mathcal{G}^{(0)}$ is Hausdorff in the subspace topology inherited from $\mathcal{G}$, and the maps $\mathbf{d}$ and $\mathbf{r}$ are local homeomorphisms. We remind the reader that étale groupoids are not required to be Hausdorff. It is a standard fact that an étale groupoid $\mathcal{G}$ is Hausdorff if and only if $\mathcal{G}^{(0)}$ is closed.
Now let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a locally compact Hausdorff space. We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [EXEL] for details). Set
$\Omega=\{(s,x):x\in D_{s^{*}s}\}\subseteq\mathcal{S}\times\mathcal{X}.$
We define an equivalence relation on $\Omega$ by
$(s,x)\sim(t,y)\Leftrightarrow x=y\text{ and }\exists e\in E_{\mathcal{S}}\text{ with }x\in D_{e}\text{ and }se=te.$
One defines a groupoid structure on $\mathcal{G}:=\Omega/{\sim}$ as follows:
  1. (1)
    $\mathcal{G}^{(2)}=\{([s,x],[t,y])\in\mathcal{G}\times\mathcal{G}:x=\alpha_{t}(y)\}$;
  2. (2)
    $s,\alpha_{t}(x)][t,x]=[st,x]$;
  3. (3)
    $s,x]^{-1}=[s^{*},\alpha_{s}(x)]$.
Equipped with the quotient topology, $\mathcal{G}$ becomes an étale groupoid, with basic open neighbourhoods of the form $\Theta(s,U)=\{[s,x]:x\in U\}$ (for $s\in\mathcal{S}$ and open $U\subseteq\mathcal{X}$). We write $\mathcal{G}=\mathcal{S}\ltimes_{\alpha}\mathcal{X}$, and call $\mathcal{G}$ the transformation groupoid (or groupoid of germs) of the action $\mathcal{S}\curvearrowright_{\alpha}\mathcal{X}$.
Every étale groupoid $\mathcal{G}$ generates a reduced C-algebra, $\textup{C}^{*}_{r}(\mathcal{G})$. As we will not have need to directly work with this C-algebra, we refer the reader to [CEPSS] for details. The relation between the two C-algebra constructions described in this section is the content of the next theorem, which is an immediate consequence of ([bussexel], Theorem 4.11).
Theorem 2.3.
Let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a compact Hausdorff space. The reduced crossed product $C_{0}(\mathcal{X})\rtimes_{r}\mathcal{S}$ is isomorphic to the reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes_{\alpha}\mathcal{X})$.
3. The Stone-Čech transformation groupoid of an inverse semigroup
Throughout this section, we identify $\beta\mathcal{S}$ with the space of ultrafilters on the discrete space $\mathcal{S}$. Given $A\subseteq\mathcal{S}$ we let
$U_{A}:=\{\xi\in\beta\mathcal{S}:A\in\xi\}.$
Sets of this form generate the topology on $\beta\mathcal{S}$. Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$; so the canonical embedding described in Example 2.2 is given by $\iota(s)=\{s\}^{\subseteq}$. In what follows we will drop the $\iota$ and consider $\mathcal{S}$ as a subset of $\beta\mathcal{S}$. We then have that $\overline{\mathcal{S}}=\beta\mathcal{S}$. Furthermore, it is a straightforward exercise to show that
$U_{A}=\overline{A},$
that is, the basic open set $U_{A}$ is the closure of $A$ in $\beta\mathcal{S}$.
Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\overline{A}\to\beta A$. With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
Theorem 3.1.
The reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\beta\mathcal{S})$ is isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
Proof.
[…]
In what remains of this paper we study the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$. In [EP16], Exel and Pardo proved many results about such actions and groupoids, so we recall some terminology and facts from there.
For $s\in\mathcal{S}$ we let
$\mathcal{J}_{s}$ $:=\{e\in E_{\mathcal{S}}:se=e\},$
$F_{s}$ $:=\{x\in\overline{s^{*}\mathcal{S}}:\lambda_{s}(x)=x\},$
$TF_{s}$ $:=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}.$
Elements of $F_{s}$ are the fixed points of $s$ and the elements of $TF_{s}$ are the trivially fixed points of $s$; this is the terminology given in [EP16].
An element $[s,x]$ is in the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ if and only if $[s,x]=[e,x]$ for some idempotent $e$, so we identify the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ with $\beta\mathcal{S}$ via $e,x]\mapsto x$. If $[s,x]=[e,x]$ for $e\in E_{\mathcal{S}}$, there is an idempotent $f$ with $x\in\overline{f\mathcal{S}}$ with $sf=ef$. Thus $sfe=efe=fe$ and so $fe\in\mathcal{J}_{s}$. We also have $[s,x]=[e,x]=[ef,x]$.
It is clear that $x\in F_{s}$ if and only if $\lambda_{s}(x)=\mathbf{r}([s,x])=\mathbf{d}([s,x])=x$. In summary, we have observed the following lemma:
Lemma 3.2.
Let $\mathcal{S}$ be an inverse semigroup. If $s\in\mathcal{S}$, then
  1. (1)
    $x\in F_{s}\iff[s,x]\in\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$;
  2. (2)
    $x\in TF_{s}\iff[s,x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$;
  3. (3)
    $TF_{s}\subseteq F_{s}$.
A key dynamical fact for our purposes is the following consequence of Frolík:
Lemma 3.3.
Let $s\in\mathcal{S}$. If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a subset $A\subseteq s^{*}\mathcal{S}$ such that $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$. In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is open.
Proof.
The result is plain if $\xi$ is a principal ultrafilter, so we may assume that $\xi$ is a free ultrafilter (and therefore that $s^{*}\mathcal{S}$ is infinite.) Let $B\subset s^{*}\mathcal{S}$ be an infinite subset with infinite complement. Either $B\in\xi$ or $s^{*}\mathcal{S}\backslash B\in\xi$; without loss of generality, take $B\in\xi$. Because $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ is a bijection, $sB\subset s\mathcal{S}$ is also an infinite subset with infinite complement. Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to s^{*}B$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$. By Frolík’s theorem ([Frolik1971], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$. The result follows after taking $A=B\cap C$. ∎
We can now derive a nice form for the fixed point set of a given $s\in\mathcal{S}$.
Lemma 3.4.
If $\mathcal{S}$ is an inverse semigroup and $s\in\mathcal{S}$, then
(1) $F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}.$
Proof.
Suppose that there is some $\xi\in\overline{s^{*}\mathcal{S}}$ such that $\lambda_{s}\xi=\xi$. By Lemma 3.3, there exists $A\subseteq s^{*}\mathcal{S}$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$. In particular, we have $sa=a$, for all $a\in A$, so $saa^{*}=aa^{*}$ for all $a\in A$. Said differently, $aa^{*}\in\mathcal{J}_{s}$, for all $a\in A$. Thus,
$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}}e\mathcal{S}\in\xi,$
because $\xi$ is a filter and filters are upwards closed. Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$. Reverse containment follows from the continuity of $\lambda_{s}$, completing the proof. ∎
Definition 3.5.
Let $\mathcal{G}$ be a groupoid. We say that
  1. (1)
    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $s(g)=r(g)$ implies $g\in\mathcal{G}^{(0)}$,
  2. (2)
[…]
  3. (3)
    $\mathcal{G}$ is essentially principal if $\{x\in\mathcal{G}^{(0)}:\mathcal{G}_{x}^{x}=\{x\}\}$ is dense in $\mathcal{G}^{(0)}$.
Proposition 3.6.
Let $s\in\mathcal{S}$ and suppose $s,x]\in\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ for some principal filter $x\in s^{*}\mathcal{S}$. Then $s,x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is always essentially principal.
Proof.
We have that $sx=x$ and so $sxx^{*}=xx^{*}$. Since $x\in xS=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in x$. Thus $s,x]=[sxx^{*},x]=[xx^{*},x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. Because principal filters are dense in $(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$, $\mathcal{S}\ltimes\beta\mathcal{S}$ is essentially principal. ∎
Theorem 3.7.
For an inverse semigroup $\mathcal{S}$, the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal;
  3. (3)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is effective.
Proof.
$(1)\implies(2)$ By [EP16], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$. Lemma 3.4 says that $F_{s}$ is the smallest closed set that contains $e\mathcal{S}$ for all $e\in\mathcal{J}_{s}$, so $F_{s}\subseteq TF_{s}$, and hence $F_{s}=TF_{s}$. Because every fixed point is trivially fixed, it follows from Lemma 3.2 that $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal.
$(2)\implies(1)$ if $\mathcal{G}$ is not Hausdorff, we can find a net $(\gamma_{i})$ in $\mathcal{G}$ with two distinct limits $\gamma,\gamma^{\prime}$. Because the range and source maps are continuous, we have $r(\gamma)=r(\gamma^{\prime})$ and $s(\gamma)=s(\gamma^{\prime})$. Because $\gamma\neq\gamma^{\prime}$, $\mathcal{G}$ is not principal, and so by the contrapositive we are done.
$(2)\iff(3)$ follows from Lemma 3.3. ∎
In light of Theorem 3.7, the key structural question one asks is when is $\mathcal{S}\ltimes\beta\mathcal{S}$ Hausdorff? Below is a criterion that produces examples.
Theorem 3.8.
Let $\mathcal{S}$ be an inverse semigroup. Then the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    for every $s\in\mathcal{S}$, there exists a finite set $F\subseteq\mathcal{J}_{s}$ such that $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$;
  3. (3)
    for every $s\in\mathcal{S}$, there exists a finite set $F\subseteq\mathcal{J}_{s}$ such that $\mathcal{J}_{s}=F^{\geqslant}$.
Proof.
$(1)\implies(2)$ By Theorem 3.7, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then $TF_{s}=F_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ for all $s\in\mathcal{S}$. This set is compact and each $\overline{e\mathcal{S}}$ is open, so there is a finite $F\subseteq\mathcal{J}_{s}$ such that
$\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}},$
where the last equality follows as $F$ is finite. Because elements of $\mathcal{S}$ are isolated in $\beta\mathcal{S}$, this implies $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
$(2)\implies(3)$ For each $e\in\mathcal{J}_{s}$, there exists $f\in F$ such that $e\in f\mathcal{S}$ so we can write $e=ft$ for some $t\in\mathcal{S}$. Thus $fe=fft=ft=e$, implying $e\leqslant f$. Conversely, if $e\leqslant f\in F\subseteq\mathcal{J}_{s}$, $se=sfe=fe=e$ and so $e\in\mathcal{J}_{s}$.
$(3)\implies(2)$ If $e\in\mathcal{J}_{s}$, then by assumption, $e\leqslant f$ for some $f\in F$; so $e\in f\mathcal{S}$. Thus, we have $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
$(2)\implies(1)$ We have
$F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}\subseteq\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=TF_{s},$
implying $F_{s}=TF_{s}$. By Theorem 3.7, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
Two elements $s,t\in\mathcal{S}$ are called compatible if both $s^{*}t$ and $st^{*}$ are idempotents. Notice that idempotents are always compatible. The inverse semigroup $\mathcal{S}$ is called complete if whenever $A\subseteq\mathcal{S}$ is a set of pairwise compatible elements, then $A$ has a least upper bound in $\mathcal{S}$ (in the natural partial order); in this case the least upper bound is denoted $\bigvee_{a\in A}a$. An inverse semigroup is called infinitely distributive if whenever $\bigvee_{a\in A}a\in\mathcal{S}$ and $s\in\mathcal{S}$, we have
$s\left(\bigvee_{a\in A}a\right)=\bigvee_{a\in A}sa\hskip 28.45274pt\text{and}\hskip 28.45274pt\left(\bigvee_{a\in A}a\right)s=\bigvee_{a\in A}as.$
An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [lawson]). The symmetric inverse monoid $\mathcal{I}(X)$ is an abstract pseudogroup. An inverse semigroup with zero $\mathcal{S}$ is $E^{*}$-unitary if $s\in\mathcal{S}$ and $\mathcal{J}_{s}\neq\{0\}$ implies $s\in E_{S}$. Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [L]. The next result is an immediate consequence of Theorem 3.8.
Corollary 3.9.
If $\mathcal{S}$ is an abstract pseudogroup or an $E^{*}$-unitary inverse semigroup, then $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff.
Proof.
If $\mathcal{S}$ is an abstract pseudogroup and $s\in\mathcal{S}$, then $\mathcal{J}_{s}$ is pairwise compatible and hence $e:=\bigvee_{a\in\mathcal{J}_{s}}a\in\mathcal{S}$. We also have that $se=\bigvee_{a\in\mathcal{J}_{s}}sa=\bigvee_{a\in\mathcal{J}_{s}}a=e$ by distributivity, so $e\in\mathcal{J}_{s}$. The finite set $F=\{e\}$ satisfies Theorem 3.8(3).
Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [EP16]. Alternatively, observe that if $\mathcal{J}_{s}$ is nonzero, then $s\in\mathcal{J}_{s}$, and thus $\mathcal{J}_{s}=\{s\}^{\geqslant}$. ∎
The main result of [starling] says that the action of $\mathcal{S}$ on its Exel spectrum is amenable if and only if every action of $\mathcal{S}$ is amenable. Our last result has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
Definition 3.10.
Say that an action $(\mathcal{S},\mathcal{X},\alpha)$ is clopen if $D^{\alpha}_{e}$ is clopen for each $e\in E_{\mathcal{S}}$.
Corollary 3.11.
For an inverse semigroup $\mathcal{S}$, the following conditions are equivalent:
  1. (1)
    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
  2. (2)
    $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff for every clopen action $(\mathcal{S},\mathcal{X},\alpha)$.
Proof.
[…]
Let $s\in\mathcal{S}$. By assumption, $\alpha_{s}:D^{\alpha}_{s^{*}s}\to D^{\alpha}_{ss^{*}}$ a homeomorphism between clopen sets. Writing $TF^{\alpha}_{s}$ for the trivially fixed points for the action $\alpha$, we have
$TF^{\alpha}_{s}=\bigcup_{e\in\mathcal{J}_{s}}D^{\alpha}_{e}.$
Applying Theorem 3.8, we can find a finite set $F\subseteq\mathcal{J}_{s}$ such that, for all $e\in\mathcal{J}_{s}$, there exists $f\in F$ with $e\leqslant f$. Since $D_{e}^{\alpha}\subseteq D_{f}^{\alpha}$ whenever $e\leqslant f$, we have
$TF^{\alpha}_{s}=\bigcup_{e\in\mathcal{J}_{s}}D^{\alpha}_{e}=\bigcup_{f\in F}D^{\alpha}_{f}.$
By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [EP16, Theorem 3.15], $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎
Corollary 3.12.
If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then Exel’s tight groupoid of $\mathcal{S}$, Exel’s universal groupoid of $\mathcal{S}$, and Paterson’s universal groupoid of $\mathcal{S}$ are all Hausdorff.
Negating Corollary 3.11 then shows that we can find many examples where $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff (and hence not principal), namely any example where the tight groupoid is not Hausdorff. See [CEPSS, Example 5.6] for an example arising from the Grigorchuk group.
Question: does Corollary 3.11 hold for non-clopen actions?

### v2

@@
 We initiate the study of the Stone-Čech transformation groupoid $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ of an inverse semigroup $\mathcal{S}$.
-This groupoid occurs as a model for the inverse semigroup Roe algebra $\mathcal{R}_{\mathcal{S}}$ of Lledo and Martinez; in particular, the action of $\mathcal{S}$ on $\beta\mathcal{S}$ is amenable in the sense of Exel and Starling if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
 We prove that the properties of being Hausdorff, principal, and effective are all equivalent for $\mathcal{G}$, and give an algebraic condition on $\mathcal{S}$ equivalent to the Hausdorffness of $\mathcal{G}$.
-Finally, we show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$.
+We show that the Hausdorffness of Exel’s tight groupoid $\mathcal{G}_{\text{tight}}(\mathcal{S})$ is necessary for the Hausdorffness of $\mathcal{G}$.
+Finally, we clarify the connection between several crossed product constructions involving this groupoid, and show that when $\mathcal{G}$ is Hausdorff and $\mathcal{S}$ has the Property (FL) of Lledó and Martínez, then $\mathcal{G}$ is amenable if and only if the reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
 1. Introduction
@@
 A semigroup $\mathcal{S}$ is called inverse if, for every $s\in\mathcal{S}$, there is a unique $s^{*}\in\mathcal{S}$ with $ss^{*}s=s$ and $s^{*}ss^{*}=s^{*}$.
-The dynamics and geometry of inverse semigroups have been productively studied in a way that generalizes the discrete group setting.
-In [lledomartinez], Lledo and Martinez described the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ of a countable inverse semigroup $\mathcal{S}$, and showed that this C-algebra was naturally isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
-In this note, we describe an étale groupoid model for $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
-Like the group case, this étale groupoid arises from a left-translation action of $\mathcal{S}$ on its Stone-Čech compactification (Theorem 3.1).
-It follows immediately that Exel’s reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable as an étale groupoid.
-Said differently, the following generalization of the discrete group case ([ozawa], Theorem 3) holds:
-Theorem 1.
-Exel’s inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable in the sense of Exel and Starling [starling].
-It is well-known that the action of a discrete group $\Gamma$ on $\beta\Gamma$ is free (see for example [NR14, Example 2.14].
-In the inverse semigroup setting, one may therefore anticipate that $\mathcal{G}=\mathcal{S}\ltimes\beta\mathcal{S}$ is always principal.
-Using a finite cover criterion of Exel and Pardo [EP16], we obtain the following generalization of this result.
+There have been recent efforts to understand the dynamics and representation theory of inverse semigroups in a way that generalizes the discrete group setting.
+The first to investigate the exactness of $\textup{C}^{*}_{r}(\mathcal{S})$ seems to be Anantharaman-Delaroche in the monograph [2], where groupoid C-algebra techniques are applied to show that $\textup{C}^{*}_{r}(\mathcal{S})\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ is exact whenever the maximum group homomorphic image $\mathcal{S}{/}\sigma$ is an exact group.
+In [14], Lledó and Martínez construct and study the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ of a countable inverse semigroup $\mathcal{S}$.
+Under a mild combinatorial assumption (admitting a finite labelling), they show that the reduced inverse semigroup C-algebra $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{R}_{\mathcal{S}}$ is nuclear.
+In this note, we consider the contravariant action $\mathcal{S}\curvearrowright\beta\mathcal{S}$, and investigate the properties of the resulting action groupoid $\mathcal{S}\ltimes\mathcal{\beta}\mathcal{S}$ and reduced C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\mathcal{\beta}\mathcal{S})\cong\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
+We first consider the question of effectiveness and principality for $\mathcal{S}\ltimes\beta\mathcal{S}$.
+These conditions always hold in the group case, and correspond to the fact that the action $\Gamma\curvearrowright\beta\Gamma$ is always free (see for example [15]*Example 2.14).
+In contrast to the group case, and somewhat surprisingly, there are inverse semigroups $\mathcal{S}$ such that $\mathcal{S}\ltimes\beta\mathcal{S}$ is not effective nor principal.
+Using a finite cover criterion of Exel and Pardo [7], we obtain the following generalization of the group case:
 Theorem 1.
@@
   1. (1)
-    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff,
+    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
   2. (2)
-    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal,
+    $\mathcal{S}\ltimes\beta\mathcal{S}$ is effective;
+  3. (3)
+    $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal;
   4. (4)
@@
 This generalizes the group case because $\mathcal{S}\ltimes\beta\mathcal{S}$ is always Hausdorff when $\mathcal{S}$ is a group.
-Using Theorem 2, we give a criterion on $\mathcal{S}$ (Theorem 3.8) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.9).
+Using Theorem 1, we give an algebraic criterion on $\mathcal{S}$ (Theorem 3.7) which implies $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff for the class of abstract pseudogroups (Corollary 3.8).
 Finally, since there are many examples of inverse semigroups with non-Hausdorff tight groupoids, condition (3) of Theorem 1 provides many examples of inverse semigroups for which $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff nor principal.
+To consider the amenability of the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ in the sense of Exel and Starling [8], we clarify the relationship between the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$ and the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ in the sense of [4].
+In [14], the uniform Roe algebra $\mathcal{R}_{\mathcal{S}}\subset\mathcal{B}(\ell^{2}(\mathcal{S}))$ is shown to embed into $\mathcal{B}(\ell^{2}(\mathcal{S})\otimes\ell^{2}(\mathcal{S}))$ as the image of a covariant representation of the action $\mathcal{S}\curvearrowright\ell^{\infty}(\mathcal{S})$.
+It follows from Theorem 1 that this C-algebra, denoted in [14] by $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$” and in this note by $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$, is not to be confused (in full generality) with the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ in the sense of [4].
+However, these C-algebras often coincide:
+Theorem 2.
+The following conditions are equivalent for an inverse semigroup $\mathcal{S}$:
+  1. (1)
+    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
+  2. (2)
+    $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}\cong\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ canonically.
+As a consequence, we obtain a generalization of a discrete group result.
+Theorem 3.
+Suppose $\mathcal{S}$ is a countable inverse semigroup with the property (FL) of Lledó and Martínez.
+If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable (in the sense of [8]) if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
+We would like to thank the anonymous referee for several suggestions and corrections from the previous version to this manuscript, and particularly for pointing out how our results in that version suggest the content of Theorem 2.
 2. Preliminaries
 In what follows, $\mathcal{S}$ denotes a countable inverse semigroup.
+We refer the reader to Lawson’s book [13] for a detailed treatment of these semigroups.
+A subset $A\subset\mathcal{S}$ is called symmetric if $s^{*}\in A$ whenever $s\in A$.
 An element $e\in\mathcal{S}$ is called idempotent if $e^{2}=e$.
@@
 It is a pleasant fact that $E_{\mathcal{S}}$ is a commutative subsemigroup of $\mathcal{S}$, and that $\mathcal{S}$ is a group if and only if $E_{\mathcal{S}}$ is a singleton.
-If $X$ is a set, then the collection of all partial bijections between subsets of $X$, denoted by $\mathcal{I}(X)$, is an inverse semigroup under the operation of partial composition called the symmetric inverse monoid on $X$.
+If $X$ is a set, then the collection $\mathcal{I}(X)$ of all partial bijections between subsets of $X$ is an inverse semigroup under the operation of partial composition.
+The semigroup $\mathcal{I}(X)$ is called the symmetric inverse monoid on $X$.
 Inverse semigroups carry a natural partial order: $s\leqslant t$ if and only if $ts^{*}s=s$. Restricted to $E_{\mathcal{S}}$, this order is given by $e\leqslant f$ if and only if $ef=e$.
-For any poset $(P,\prec)$ and $A\subseteq P$ we will use the notation
+For any poset $(P,\prec)$ and subset $A\subseteq P$ we will use the notation
 $A^{\prec}:=\{b\in P:a\prec b\text{ for some }a\in A\}.$
 In this paper, we only consider the partial orders $\subseteq$ on the power set of $\mathcal{S}$, and $\leqslant$ or $\geqslant$ on $\mathcal{S}$.
-We now introduce actions of inverse semigroups on topological spaces and C-algebras.
-Although this note is only interested in two such actions (one on $\beta\mathcal{S}$, and one on $\ell^{\infty}(\mathcal{S})$), we feel it will be beneficial to introduce the theory in somewhat excessive generality.
-Let $\mathcal{A}$ be a C-algebra.
-An action of an inverse semigroup $\mathcal{S}$ on $\mathcal{A}$ is a semigroup homomorphism $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{A})$ such that, for each $s\in\mathcal{S}$,
-  1. (1)
-    The domain and codomain of $\alpha_{s}$ is a closed, two-sided ideal in $\mathcal{A}$;
-  2. (2)
-    The partial bijection $\alpha_{s}$ is a *-homomorphism.
-Given an idempotent $e\in E_{\mathcal{S}}$, we set $I_{e}=\text{dom}(\alpha_{e})$.
-We observe that $\alpha_{s}$ is a *-isomorphism of $I_{s^{*}s}$ onto $I_{ss^{*}}$.
-The next example, central in [lledomartinez], is also the main C-action we will consider in this note.
-Example 2.1.
-Let $\mathcal{A}=\ell^{\infty}(\mathcal{S})$.
-For each $e\in E_{\mathcal{S}}$, define $I_{e}=\{f\in\mathcal{A}:\text{supp}(f)\subseteq e\mathcal{S}\}$.
-For each $s\in\mathcal{S}$, define $\alpha_{s}:I_{s^{*}s}\to I_{ss^{*}}$ by $\alpha_{s}f](t)=f(s^{*}t)$.
-The mapping $s\in\mathcal{S}\mapsto\alpha_{s}\in\mathcal{I}(\mathcal{S})$ is an action of $\mathcal{S}$ on $\mathcal{A}$.
-The C-algebra $\ell^{\infty}(\mathcal{S})$ is *-isomorphic to $C(\beta\mathcal{S})$, where $\beta\mathcal{S}$ is the Stone-Čech compactification of $\mathcal{S}$.
-At the level of compact Hausdorff spaces, inverse semigroup actions are described as follows (see [EXEL] for details): an action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
+An action of $\mathcal{S}$ on a compact Hausdorff space $\mathcal{X}$ is a triple $(\mathcal{S},\mathcal{X},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{X})$ is a semigroup homomorphism such that, for all $s\in\mathcal{S}$,
   1. (1)
@@
     $\text{dom}(\alpha_{s})$ is open, for all $s\in\mathcal{S}$.
-[…]
-Example 2.2 ([gondek24], Section 3.2).
-Let us define an inverse semigroup action of $\mathcal{S}$ on the Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding.
+One can see that the domain of $\alpha_{s}$ is the same as that of $\alpha_{s^{*}s}$.
+For each $e\in E_{\mathcal{S}}$ we write $D_{e}^{\alpha}$ for the domain of $\alpha_{e}$ (or simply $D_{e}$ if the action is understood), so that $\alpha_{s}:D_{s^{*}s}^{\alpha}\to D_{ss^{*}}^{\alpha}$.
+The main action we are concerned with in this paper is described below.
+Example 2.1 ([11], Section 3.2).
+We define the canonical action of an inverse semigroup $\mathcal{S}$ on its Stone-Čech compactification $(\beta\mathcal{S},\iota)$, where $\mathcal{S}$ is viewed as a discrete space and $\iota:\mathcal{S}\to\beta\mathcal{S}$ is the canonical embedding.
 For each $s\in\mathcal{S}$, define a bijection $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ by $\lambda_{s}t=st$ ($t\in s^{*}\mathcal{S}$).
 By the universal property of $\beta\mathcal{S}$, $\lambda_{s}$ extends to a homeomorphism $\lambda_{s}:\overline{s^{*}\mathcal{S}}\to\overline{s\mathcal{S}}$, where $\overline{s^{*}\mathcal{S}}:=\overline{\iota(s^{*}\mathcal{S})}^{\beta}\subseteq\beta\mathcal{S}$.
-We observe that the domain and codomain of $\lambda_{s}$ are always a compact, open subsets of $\beta\mathcal{S}$.
+We observe that the domain and codomain of $\lambda_{s}$ are always compact, open subsets of $\beta\mathcal{S}$.
 It is routine to check that the mapping $\mathcal{S}\ni s\mapsto\lambda_{s}\in\mathcal{I}(\beta\mathcal{S})$ is a semigroup homomorphism, and therefore an action, which we shall henceforth refer to by $\mathcal{S}\curvearrowright\beta\mathcal{S}$.
-The contravariant action $\hat{\lambda}$ of $\mathcal{S}$ on $C(\beta\mathcal{S})$ is the action in Example 2.1, after applying the identification $j:C(\beta\mathcal{S})\ni f\mapsto f\circ\iota\in\ell^{\infty}(\mathcal{S})$.
-Indeed, given $f\in I_{s^{*}s}\subseteq\ell^{\infty}(\mathcal{S})$, we have $j(f)](t)=f(\iota(t))=0\Leftrightarrow\iota(t)\not\in\overline{s^{*}\mathcal{S}}\Leftrightarrow t\not\in s^{*}\mathcal{S}$, so that $j(\text{dom}(\hat{\lambda}_{s}))=I_{s^{*}s}$, and
-$[\alpha_{s}(j(f))](t)=[j(f)](s^{*}t)=f(\iota(s^{*}t))=f(\lambda_{s^{*}}\iota(t))=[\hat{\lambda}_{s}f](t),$
-for all $f\in I_{s^{*}s}$.
-Given a C-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ and an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{A}$, there is a standard construction for a reduced crossed product C-algebra $\mathcal{A}\rtimes_{r}\mathcal{S}\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$, which does not depend on the inclusion $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$ (for details, see [EXEL]).
-We will not work directly with this C-algebra, so we only sketch the construction.
-For each $s\in\mathcal{S}$, we define the maps $V_{s}:\mathcal{S}\to\mathcal{B}(\ell^{2}(\mathcal{S}))$ and $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}\otimes\mathcal{\ell}^{2}(\mathcal{S}))$ by
-$V_{s}\delta_{t}=\begin{cases}\delta_{st}&t\in s^{*}\mathcal{S}\\ 0&\text{otherwise}\end{cases}\qquad$ $\text{and}\qquad\pi(a)(\xi\otimes\delta_{t})=\begin{cases}\alpha_{t^{*}}(a)\xi\otimes\delta_{t}&t\in I_{s^{*}s}\\ 0&\text{otherwise}.\end{cases}$
-The reduced crossed product is the C-subalgebra of $\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S}))$ defined by
-$\mathcal{A}\rtimes_{r}\mathcal{S}:=\textup{C}^{*}(\{1\otimes V_{s}:s\in\mathcal{S}\}\cup\pi(A))\subseteq\mathcal{B}(\mathcal{H}\otimes\ell^{2}(\mathcal{S})).$
-We are interested in describing an étale groupoid $\mathcal{G}$ with the property that $\text{C}^{*}_{r}(\mathcal{G})=\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
-In preparation for this, let us establish conventions for étale groupoids in this note.
+The appropriate way to interpret the dynamics of an inverse semigroup action $\mathcal{S}\curvearrowright\mathcal{X}$ is through the corresponding transformation groupoid, denoted by $\mathcal{S}\ltimes\mathcal{X}$.
+Before reminding the reader of this construction, we establish our conventions for étale groupoids.
+We direct the reader to the lecture notes [21] of Sims for details.
+A groupoid is a small category in which every arrow is invertible.
 Given a groupoid $\mathcal{G}$, we will denote by $\mathcal{G}^{(2)}\subset\mathcal{G}\times\mathcal{G}$ the composable pairs, and $\mathcal{G}^{(0)}=\{\gamma^{-1}\gamma:\gamma\in\mathcal{G}\}$ the unit space.
@@
 A topological groupoid is called étale if it is locally compact, $\mathcal{G}^{(0)}$ is Hausdorff in the subspace topology inherited from $\mathcal{G}$, and the maps $\mathbf{d}$ and $\mathbf{r}$ are local homeomorphisms.
+A bisection is a subset $U\subset\mathcal{G}$ such that $\mathbf{d}|_{U}$ and $\mathbf{r}|_{U}$ are local homeomorphisms.
 We remind the reader that étale groupoids are not required to be Hausdorff.
@@
 Now let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a locally compact Hausdorff space.
-We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [EXEL] for details).
+We associate an étale groupoid to $(\mathcal{S},\mathcal{X},\alpha)$ as follows (see [9] for details).
 Set
@@
 We write $\mathcal{G}=\mathcal{S}\ltimes_{\alpha}\mathcal{X}$, and call $\mathcal{G}$ the transformation groupoid (or groupoid of germs) of the action $\mathcal{S}\curvearrowright_{\alpha}\mathcal{X}$.
-Every étale groupoid $\mathcal{G}$ generates a reduced C-algebra, $\textup{C}^{*}_{r}(\mathcal{G})$.
-As we will not have need to directly work with this C-algebra, we refer the reader to [CEPSS] for details.
-The relation between the two C-algebra constructions described in this section is the content of the next theorem, which is an immediate consequence of ([bussexel], Theorem 4.11).
-Theorem 2.3.
-Let $(\mathcal{S},\mathcal{X},\alpha)$ be an inverse semigroup action, where $\mathcal{X}$ is a compact Hausdorff space.
-The reduced crossed product $C_{0}(\mathcal{X})\rtimes_{r}\mathcal{S}$ is isomorphic to the reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes_{\alpha}\mathcal{X})$.
 3. The Stone-Čech transformation groupoid of an inverse semigroup
@@
 Sets of this form generate the topology on $\beta\mathcal{S}$.
-Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$; so the canonical embedding described in Example 2.2 is given by $\iota(s)=\{s\}^{\subseteq}$.
+Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$, so the canonical embedding described in Example 2.1 is given by $\iota(s)=\{s\}^{\subseteq}$.
 In what follows we will drop the $\iota$ and consider $\mathcal{S}$ as a subset of $\beta\mathcal{S}$.
 We then have that $\overline{\mathcal{S}}=\beta\mathcal{S}$.
-Furthermore, it is a straightforward exercise to show that
-$U_{A}=\overline{A},$
+Now observe that
+$U_{A}=\overline{A};$
 that is, the basic open set $U_{A}$ is the closure of $A$ in $\beta\mathcal{S}$.
-Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\overline{A}\to\beta A$. With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
-Theorem 3.1.
-The reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\beta\mathcal{S})$ is isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.
-In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
-Proof.
-[…]
-In what remains of this paper we study the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$.
-In [EP16], Exel and Pardo proved many results about such actions and groupoids, so we recall some terminology and facts from there.
+To see this, note that $\xi\in\overline{A}$ if and only if, for all $B\in\xi$, there is some $a\in A$ with $a\in B$.
+This condition forces $A\in\xi$, because $\xi$ is an ultrafilter and otherwise one would have $\mathcal{S}\backslash A\in\xi$.
+Conversely, if $A\in\xi$, then already $A\cap B\neq\emptyset$ for all $B\in\xi$, so $\xi\in\overline{A}$.
+Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\beta\mathcal{S}\supset\overline{A}\to\beta A$, for all $A\subset\mathcal{S}$.
+With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
+We now turn to studying the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$.
+In [7], Exel and Pardo proved many general results about such actions and groupoids, so we recall some terminology and facts from there.
 For $s\in\mathcal{S}$ we let
@@
 $TF_{s}$ $:=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}.$
-Elements of $F_{s}$ are the fixed points of $s$ and the elements of $TF_{s}$ are the trivially fixed points of $s$; this is the terminology given in [EP16].
+Following the terminology given in [7], the elements of $F_{s}$ are called the fixed points of $s$, and the elements of $TF_{s}$ are called the trivially fixed points of $s$.
 An element $[s,x]$ is in the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ if and only if $[s,x]=[e,x]$ for some idempotent $e$, so we identify the unit space of $\mathcal{S}\ltimes\beta\mathcal{S}$ with $\beta\mathcal{S}$ via $e,x]\mapsto x$.
@@
     $TF_{s}\subseteq F_{s}$.
+Consequently, $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal if and only if $TF_{s}=F_{s}$ for all $s\in\mathcal{S}$.
 A key dynamical fact for our purposes is the following consequence of Frolík:
@@
 Let $s\in\mathcal{S}$.
-If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a subset $A\subseteq s^{*}\mathcal{S}$ such that $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$.
-In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is open.
+If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a nonempty subset $A\subseteq s^{*}\mathcal{S}$ such that $A\in\xi$ and $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$.
+In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is nonempty and open.
 Proof.
@@
 Because $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ is a bijection, $sB\subset s\mathcal{S}$ is also an infinite subset with infinite complement.
-Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to s^{*}B$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$.
-By Frolík’s theorem ([Frolik1971], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$.
+Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to sB$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$.
+By Frolík’s theorem ([10], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$.
 The result follows after taking $A=B\cap C$.
@@
 Suppose that there is some $\xi\in\overline{s^{*}\mathcal{S}}$ such that $\lambda_{s}\xi=\xi$.
-By Lemma 3.3, there exists $A\subseteq s^{*}\mathcal{S}$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$.
+By Lemma 3.2, there exists $A\subseteq s^{*}\mathcal{S}$ with $A\in\xi$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$.
 In particular, we have $sa=a$, for all $a\in A$, so $saa^{*}=aa^{*}$ for all $a\in A$.
@@
 Thus,
-$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}}e\mathcal{S}\in\xi,$
+$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}\in\xi,$
 because $\xi$ is a filter and filters are upwards closed.
-Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$.
+Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$.
 Reverse containment follows from the continuity of $\lambda_{s}$, completing the proof.
@@
 Definition 3.4.
-Let $\mathcal{G}$ be a groupoid. We say that
+Let $\mathcal{G}$ be a topological groupoid. We say that
   1. (1)
-    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $s(g)=r(g)$ implies $g\in\mathcal{G}^{(0)}$,
+    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $\mathbf{d}(g)=\mathbf{r}(g)$ implies $g\in\mathcal{G}^{(0)}$,
   2. (2)
@@
     $\mathcal{G}$ is essentially principal if $\{x\in\mathcal{G}^{(0)}:\mathcal{G}_{x}^{x}=\{x\}\}$ is dense in $\mathcal{G}^{(0)}$.
+We point out to the reader that some authors use the term topologically free to mean what we have defined here as essentially principal.
 Proposition 3.5.
@@
 We have that $sx=x$ and so $sxx^{*}=xx^{*}$.
-Since $x\in xS=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in x$.
+Since $x\in x\mathcal{S}=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in\{x\}^{\subseteq}$.
 Thus $s,x]=[sxx^{*},x]=[xx^{*},x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$.
@@
 Proof.
-$(1)\implies(2)$ By [EP16], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$.
+$(1)\implies(2)$ By [7], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$.
+Because $\overline{s^{*}\mathcal{S}}$ is already closed, this is equivalent to $TF_{s}$ being closed.
 Lemma 3.3 says that $F_{s}$ is the smallest closed set that contains $e\mathcal{S}$ for all $e\in\mathcal{J}_{s}$, so $F_{s}\subseteq TF_{s}$, and hence $F_{s}=TF_{s}$.
@@
 Proof.
-$(1)\implies(2)$ By Theorem 3.7, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then $TF_{s}=F_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ for all $s\in\mathcal{S}$.
+$(1)\implies(2)$ By Theorem 3.6, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then it is principal and so $F_{s}=TF_{s}$ for all $s\in\mathcal{S}$.
+We always have $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ and Lemma 3.3 implies $F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$.
+Thus for all $s\in\mathcal{S}$ we have $\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$.
 This set is compact and each $\overline{e\mathcal{S}}$ is open, so there is a finite $F\subseteq\mathcal{J}_{s}$ such that
 $\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}},$
-where the last equality follows as $F$ is finite.
+where the first equality is the expression $F_{s}=TF_{s}$ and the last equality follows as $F$ is finite.
 Because elements of $\mathcal{S}$ are isolated in $\beta\mathcal{S}$, this implies $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
@@
 $F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\overline{\bigcup_{f\in F}f\mathcal{S}}=\bigcup_{f\in F}\overline{f\mathcal{S}}\subseteq\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}=TF_{s},$
-implying $F_{s}=TF_{s}$. By Theorem 3.7, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
+where the middle equality follows because the union in question is finite. This implies $F_{s}=TF_{s}$. By Theorem 3.6, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
 Two elements $s,t\in\mathcal{S}$ are called compatible if both $s^{*}t$ and $st^{*}$ are idempotents.
@@
 $s\left(\bigvee_{a\in A}a\right)=\bigvee_{a\in A}sa\hskip 28.45274pt\text{and}\hskip 28.45274pt\left(\bigvee_{a\in A}a\right)s=\bigvee_{a\in A}as.$
-An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [lawson]).
+An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [12]).
+This notion is not to be confused with other inequivalent definitions of pseudogroup, such as a pseudogroup in the sense of Renault (cf.
+[20]).
 The symmetric inverse monoid $\mathcal{I}(X)$ is an abstract pseudogroup.
 An inverse semigroup with zero $\mathcal{S}$ is $E^{*}$-unitary if $s\in\mathcal{S}$ and $\mathcal{J}_{s}\neq\{0\}$ implies $s\in E_{S}$.
-Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [L].
+Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [13].
 The next result is an immediate consequence of Theorem 3.7.
@@
 The finite set $F=\{e\}$ satisfies Theorem 3.7(3).
-Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [EP16].
+Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [7].
 Alternatively, observe that if $\mathcal{J}_{s}$ is nonzero, then $s\in\mathcal{J}_{s}$, and thus $\mathcal{J}_{s}=\{s\}^{\geqslant}$.
 ∎
-The main result of [starling] says that the action of $\mathcal{S}$ on its Exel spectrum is amenable if and only if every action of $\mathcal{S}$ is amenable.
-Our last result has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
+The main result of [8] says that the action of $\mathcal{S}$ on its spectrum $\widehat{E}_{0}$ (the space of nonzero semicharacters on $E_{\mathcal{S}}$) is amenable if and only if every action of $\mathcal{S}$ is amenable.
+Our last result in this section has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
 Definition 3.9.
@@
 $TF^{\alpha}_{s}=\bigcup_{e\in\mathcal{J}_{s}}D^{\alpha}_{e}=\bigcup_{f\in F}D^{\alpha}_{f}.$
-By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [EP16, Theorem 3.15], $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎
+By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [7]*Theorem 3.15, $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎
 Corollary 3.11.
 If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then Exel’s tight groupoid of $\mathcal{S}$, Exel’s universal groupoid of $\mathcal{S}$, and Paterson’s universal groupoid of $\mathcal{S}$ are all Hausdorff.
+For more details, Exel’s universal and tight groupoid constructions can be found in [9]; Paterson’s universal groupoid construction can be found in [18].
 Negating Corollary 3.10 then shows that we can find many examples where $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff (and hence not principal), namely any example where the tight groupoid is not Hausdorff.
-See [CEPSS, Example 5.6] for an example arising from the Grigorchuk group.
-Question: does Corollary 3.11 hold for non-clopen actions?
+See [16]*Example 5.6 for an example arising from the Grigorchuk group.
+We thank the anonymous referee for the following example, which shows that Corollary 3.10 does not hold for non-clopen actions.
+Let $\mathcal{S}=\{0,s,1\}$, where 0 and $1$ act as zero and identity elements of $\mathcal{S}$, and $s^{2}=1$.
+Since $\mathcal{S}$ is finite, $\mathcal{S}\ltimes\beta\mathcal{S}$ is discrete, and in particular Hausdorff.
+However, $\mathcal{S}$ acts trivially on $[0,1]$, with $D_{1}=D_{s}=[0,1]$ and $D_{0}=[0,1)$; the resulting action groupoid $\mathcal{S}\ltimes[0,1]$ is not Hausdorff, because the elements $[s,1]$ and $[0,1]$ cannot be separated by disjoint neighbourhoods.
+4. Amenability and the uniform Roe algebra
+In this last, self-contained section, we clarify the relationship between several crossed product constructions associated to the Stone-Čech action $\lambda:\mathcal{S}\curvearrowright\beta\mathcal{S}$, and relate the amenability of the Stone-Čech action to the nuclearity and exactness of these constructions.
+We begin by collecting several definitions and existing results.
+The uniform Roe algebra $\mathcal{R}_{\mathcal{S}}$.
+Given an inverse semigroup $\mathcal{S}$, we define the left regular representation $\lambda:\mathcal{S}\to\mathcal{B}(\ell^{2}(\mathcal{S}))$ by
+$\lambda_{s}\delta_{t}=\begin{cases}\delta_{st}&t\in s^{*}\mathcal{S};\\ 0&\text{otherwise.}\end{cases}$
+The C-algebra generated by $\{\lambda_{s}:s\in\mathcal{S}\}$ in $\mathcal{B}(\ell^{2}(\mathcal{S}))$ is called the reduced C-algebra of $\mathcal{S}$, denoted by $\textup{C}^{*}_{r}(\mathcal{S})$.
+The C-algebra generated by $\{M_{f}\lambda_{s}:s\in\mathcal{S},f\in\ell^{\infty}(\mathcal{S})\}$, where $\ell^{\infty}(\mathcal{S})\ni f\mapsto M_{f}\in\mathcal{B}(\ell^{2}(\mathcal{S}))$ is the diagonal embedding, is called the uniform Roe algebra of $\mathcal{S}$, denoted by $\mathcal{R}_{\mathcal{S}}$.
+Note that we have an inclusion $\textup{C}^{*}_{r}(\mathcal{S})\subset\mathcal{R}_{\mathcal{S}}$.
+For the convenience of the reader, we recall the following terminology from [14]: an inverse semigroup $\mathcal{S}$ has the property (FL) if there is a symmetric, generating subset $K\subset\mathcal{S}$, a natural number $N\in\mathbb{N}$, and a finite subset $K_{0}\subset K$ such that: for all $s,t\in\mathcal{S}$ with $s^{*}s=t^{*}t$ and $t\in Ks$, we have $y\in(K_{1})^{n}x$, for some $n\leq N$.
+Clearly, every finitely generated inverse semigroup has (FL); on the other hand, the free inverse semigroup on countably many generators has (FL) (see [14]*Proposition 3.18.) The relevance of this technical condition is in the following theorem, which generalizes [17]*Theorem 3:
+Theorem 4.1 ([14]*Theorem 4.18).
+If $\mathcal{S}$ is an inverse semigroup with (FL), then the following conditions are equivalent:
+  1. (1)
+    $\textup{C}^{*}_{r}(\mathcal{S})$ is exact;
+  2. (2)
+    $\mathcal{R}_{\mathcal{S}}$ is exact;
+  3. (3)
+    $\mathcal{R}_{\mathcal{S}}$ is nuclear.
+$\mathcal{R}_{\mathcal{S}}$ as the image of a covariant representation.
+An action of an inverse semigroup $\mathcal{S}$ on a unital C-algebra $\mathcal{A}$ is a triple $(\mathcal{S},\mathcal{A},\alpha)$, where $\alpha:\mathcal{S}\to\mathcal{I}(\mathcal{A})$ is a semigroup homomorphism such that, for each $s\in\mathcal{S}$,
+  1. (1)
+    The domain and codomain of $\alpha_{s}$ is a closed, two-sided ideal in $\mathcal{A}$;
+  2. (2)
+    The partial bijection $\alpha_{s}$ is a *-homomorphism.
+Given an idempotent $e\in E_{\mathcal{S}}$, we set $J_{e}=\text{dom}(\alpha_{e})$. We observe that $\alpha_{s}$ is a *-isomorphism of $J_{s^{*}s}$ onto $J_{ss^{*}}$.
+[…]
+Example 4.2 ([14]*Proposition 2.4).
+Let $\mathcal{A}=\ell^{\infty}(\mathcal{S})$.
+For each $e\in E_{\mathcal{S}}$, define $J_{e}=\{f\in\mathcal{A}:\text{supp}(f)\subseteq e\mathcal{S}\}$.
+For each $s\in\mathcal{S}$, define $\alpha_{s}:J_{s^{*}s}\to J_{ss^{*}}$ by $\alpha_{s}f](t)=f(s^{*}t)$.
+The mapping $s\in\mathcal{S}\mapsto\alpha_{s}\in\mathcal{I}(\mathcal{S})$ is an action of $\mathcal{S}$ on $\mathcal{A}$.
+Using the canonical isomorphism $C(\beta\mathcal{S})\cong\ell^{\infty}(\mathcal{S})$, the reader can easily verify that the action $\alpha:\mathcal{S}\curvearrowright\ell^{\infty}(\mathcal{S})$ is the contravariant action $\hat{\lambda}$ obtained from the Stone-Čech action $\lambda:\mathcal{S}\curvearrowright\beta\mathcal{S}$ defined in Example 2.1.
+A covariant representation of an inverse semigroup action $\alpha:\mathcal{S}\curvearrowright\mathcal{A}$ on a unital C-algebra $\mathcal{A}$ is a triple $(\mathcal{H},\pi,\sigma)$, where $\mathcal{H}$ is a Hilbert space, $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H})$ is a *-homomorphism, and $\sigma:\mathcal{S}\to\mathcal{B}(\mathcal{H})$ is a semigroup homomorphism (under composition) such that, for all $e\in E_{\mathcal{S}}$, $s\in\mathcal{S}$, and $a\in J_{s^{*}s}$,
+  1. (1)
+    $\sigma_{s^{*}}=\sigma_{s}^{*}$;
+  2. (2)
+    $\sigma_{s^{*}}\pi(a)\sigma_{s}=\pi(\alpha_{s}a)$, and
+  3. (3)
+    (Non-degeneracy) $\overline{\pi(J_{e})\mathcal{H}}=\sigma_{e}\mathcal{H}$.
+In [14]*Section 2, the following triple $(\mathcal{H},\pi,\sigma)$ is considered for the action $\mathcal{S}\curvearrowright\ell^{\infty}(\mathcal{S})$: $\mathcal{H}=\ell^{2}(\mathcal{S})\otimes\ell^{2}(\mathcal{S})$, and the maps $\pi:\ell^{\infty}(\mathcal{S})\to\mathcal{B}(\mathcal{H})$ and $\sigma:\mathcal{S}\to\mathcal{B}(\mathcal{H})$ are determined by the formulas
+$[\pi(f)](\delta_{s}\otimes\delta_{t})=\begin{cases}f(ts)\delta_{s}\otimes\delta_{t}&s\in t^{*}\mathcal{S}\\ 0&\text{otherwise}\end{cases}$
+and
+$[\sigma_{r}](\delta_{s}\otimes\delta_{t})=\begin{cases}\delta_{s}\otimes\delta_{rt}&t\in r^{*}\mathcal{S}\\ 0&\text{otherwise.}\end{cases}$
+It is easily verified that the triple $(\mathcal{H},\pi,\sigma)$ satisfies the first two conditions of a covariant representation.
+The non-degeneracy condition (3) is immediate, since $\pi(J_{e})\mathcal{H}=\pi(\ell^{\infty}(e\mathcal{S}))\mathcal{H}=\ell^{2}(\mathcal{S})\otimes\ell^{2}(e\mathcal{S})=\sigma_{e}\mathcal{H}$ (where the embeddings $\ell^{p}(e\mathcal{S})\subset\ell^{p}(\mathcal{S})$ for $p\in\{2,\infty\}$ are obtained by setting $f(t)=0$ for $t\not\in e\mathcal{S}$, $f\in\ell^{p}(e\mathcal{S})$.) Thus, the triple $(\mathcal{H},\pi,\sigma)$ forms a covariant representation for the action $\mathcal{S}\curvearrowright\ell^{\infty}(\mathcal{S})$.
+We denote the C-subalgebra generated by $(\mathcal{H},\pi,\sigma)$ by $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}\subset\mathcal{B}(\ell^{2}(\mathcal{S})\otimes\ell^{2}(\mathcal{S}))$.
+Note: in [14], the C-algebra $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$ is denoted by $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$.” We will reserve this notation for the reduced crossed product of Buss-Exel-Meyer [4], defined shortly.
+For now, we record the following result:
+Theorem 4.3 ([14]*Theorem 2.5).
+If $\mathcal{S}$ is an inverse semigroup, then the mapping $\pi(f)\sigma_{s}\mapsto M_{f}\lambda_{s}$ determines an isomorphism
+$\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}\cong\mathcal{R}_{\mathcal{S}}.$
+4.1. Reduced groupoid C-algebras
+We will shortly compare the above constructions to the reduced crossed product of Buss, Exel, and Meyer, which is defined in full generality in [4].
+Our working definition for this construction in the present note will be from a groupoid model.
+To keep this section brief, we will only summarize the key facts we need; for details on the construction of the reduced C-algebra of general étale groupoids, see [4].
+Recall that if $\mathcal{G}$ is an étale groupoid, then the *-vector space $B_{c}(\mathcal{G})$ of compactly supported Borel functions on $\mathcal{G}$ is equipped with a *-algebra structure under the convolution product
+$f*g(\gamma)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
+The set $\mathscr{C}_{c}(\mathcal{G}):=\text{span}(\bigcup_{U\in\mathscr{B}}C_{c}(U))\subset B_{c}(\mathcal{G})$, where $\mathscr{B}$ is the set of open bisections in $\mathcal{G}$, is a *-subalgebra.
+For each $x\in\mathcal{G}^{(0)}$, let $\rho_{x}:\mathscr{C}_{c}(\mathcal{G})\to\mathcal{B}(\ell^{2}(\mathcal{G}_{x}))$ denote the *-representation of $\mathscr{C}_{c}(\mathcal{G})$ given by
+$\rho_{x}(f)\delta_{\gamma}=\sum_{\alpha\in\mathcal{G}_{\mathbf{r}(x)}}f(\alpha)\delta_{\alpha\gamma},\hskip 28.45274ptf\in\mathscr{C}_{c}(\mathcal{G}),\gamma\in\mathcal{G}_{x}.$
+The reduced C-algebra of $\mathcal{G}$, denoted $\textup{C}_{r}^{*}(\mathcal{G})$, is the completion of the image of $\mathscr{C}_{c}(\mathcal{G})$ under the representation $\bigoplus_{x\in\mathcal{G}^{(0)}}\rho_{x}$.
+By [19]*Proposition 1.11, this representation is faithful.
+In particular, if $\mathcal{G}^{(0)}$ is compact, then we have a canonical inclusion $C(\mathcal{G}^{(0)})\subset\textup{C}^{*}_{r}(\mathcal{G})$.
+Let $A$ be a unital C-algebra, and $B\subset A$ a unital C-subalgebra.
+Recall that a weak conditional expectation $\mathbb{E}:A\to B$ is a unital completely positive map $\mathbb{E}:A\to B^{**}$ such that $\mathbb{E}|_{B}=\text{id}_{B}$ and $\mathbb{E}(bab^{\prime})=b\mathbb{E}(a)b^{\prime}$, for all $a\in A$, $b,b^{\prime}\in B$.
+Theorem 4.4.
+[19]*Proposition 4.8 Let $\mathcal{G}$ be an étale groupoid with compact unit space $\mathcal{G}^{(0)}$.
+  1. (1)
+    There is a faithful weak conditional expectation $\mathbb{E}:\textup{C}^{*}_{r}(\mathcal{G})\to C(\mathcal{G}^{(0)})^{**}$, which has range $C(\mathcal{G}^{(0)})$ if and only if $\mathcal{G}$ is Hausdorff.
+  2. (2)
+    If $\mathcal{G}$ is Hausdorff, then $\mathscr{C}_{c}(\mathcal{G})=C_{c}(\mathcal{G})$, and for all $f\in C_{c}(\mathcal{G})$ we have $\mathbb{E}(f)=f|_{\mathcal{G}^{(0)}}.$
+Now let $\mathcal{S}$ be a unital inverse semigroup acting on a compact Hausdorff space $\mathcal{X}$, and consider $\mathcal{G}=\mathcal{S}\ltimes\mathcal{X}$.
+Given $s\in\mathcal{S}$ and $f\in J_{ss^{*}}$, we define an element $fs\in\mathscr{C}_{c}(\mathcal{G})$ by
+$fs([t,x])=\begin{cases}f([t,x])&\text{if }[t,x]=[s,x];\\ 0&\text{otherwise.}\end{cases}$
+Theorem 4.5.
+Let $\mathcal{S}$, $\mathcal{X}$, and $\mathcal{G}$ be as above, and let $C(\mathcal{X})\rtimes_{r}\mathcal{S}$ be the reduced crossed product of $(\mathcal{S},\mathcal{X})$ in the sense of [4].
+  1. (1)
+    [9]*Proposition 3.10 The functions fsfs $(s\in\mathcal{S},f\in J_{ss^{*}})$ span $\mathscr{C}_{c}(\mathcal{G})$.
+  2. (2)
+    [9]*Theorem 8.5 If $(\mathcal{H},\pi,\sigma)$ is a covariant representation of $(\mathcal{S},\mathcal{X})$, then the formula
+    $(\pi\times\sigma)(fs)=\pi(f)\sigma_{s}$
+    determines a unique *-representation $\pi\times\sigma:\mathscr{C}_{c}(\mathcal{G})\to\mathcal{B}(\mathcal{H})$.
+  3. (3)
+    [5]*Theorem 4.11 $\textup{C}^{*}_{r}(\mathcal{G})\cong C(\mathcal{X})\rtimes_{r}\mathcal{S}$.
+Proposition 4.6.
+Let $\mathcal{S}$ be an inverse semigroup. If $\mathcal{S}\rtimes\beta\mathcal{S}$ is Hausdorff, then:
+  1. (1)
+    The *-homomorphism $\pi\times\sigma:C_{c}(\mathcal{S}\rtimes\beta\mathcal{S})\to\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$ is faithful.
+  2. (2)
+    The faithful conditional expectation $\mathbb{E}:C_{c}(\mathcal{S}\rtimes\beta\mathcal{S})\to\ell^{\infty}(\mathcal{S})$ descends to a faithful conditional expectation on $\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$.
+Proof.
+We prove (2) first. This follows from Theorem 4.5(1) because
+$[\mathbb{E}(fs)]\delta_{t}=\langle M_{f}\lambda_{s}\delta_{t},\delta_{t}\rangle,$
+for all $s,t\in\mathcal{S}$ and $f\in J_{ss^{*}}$, and because the mapping
+$\mathbb{E}^{\prime}:\mathcal{B}(\ell^{2}(\mathcal{S}))\ni T\mapsto[\mathcal{S}\ni t\mapsto\langle T\delta_{t},\delta_{t}\rangle\in\mathbb{C}]\in\ell^{\infty}(\mathcal{S})$
+is a faithful conditional expectation.
+(1) follows because if $a\in C_{c}(\mathcal{S}\rtimes\beta\mathcal{S})$, then $(\pi\times\sigma)(a)=0\Leftrightarrow(\pi\times\sigma)(a^{*}a)=0\Leftrightarrow\mathbb{E}^{\prime}((\pi\times\sigma)(a^{*}a))=0\Leftrightarrow\mathbb{E}(a^{*}a)=0\Leftrightarrow a=0$.
+∎
+Corollary 4.7.
+If $\mathcal{S}$ is an inverse semigroup, then the following conditions are equivalent:
+  1. (1)
+    $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff;
+  2. (2)
+    The map $fs\mapsto M_{f}\lambda_{s}$ extends to an isomorphism $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}\cong\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$.
+Proof.
+If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then by Proposition 4.6, both $\mathcal{A}_{1}=\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$ and $\mathcal{A}_{2}=\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ are completions of $C_{c}(\mathcal{S}\rtimes\beta\mathcal{S})$ which admit faithful extensions of the conditional expectation $\mathbb{E}$.
+It follows that $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ act faithfully on the left multiplication Hilbert $\ell^{\infty}(\mathcal{S})$-modules $L^{2}(\mathcal{A}_{1},\mathbb{E})$ and $L^{2}(\mathcal{A}_{2},\mathbb{E})$ (respectively), and that the identity map on $C_{c}(\mathcal{S}\rtimes\beta\mathcal{S})$ extends to a unitary $\mathcal{U}:L^{2}(\mathcal{A}_{1},\mathbb{E})\to L^{2}(\mathcal{A}_{2},\mathbb{E})$, which implements an isomorphism between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$.
+Conversely, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is not Hausdorff, then (by Theorem 4.4(1)) the weak conditional expectation $\mathbb{E}:\mathcal{A}_{2}\to\ell^{\infty}(\mathcal{S})^{**}$ takes values outside of $\ell^{\infty}(\mathcal{S})$, while the conditional expectation $\mathbb{E}^{\prime}:\mathcal{A}_{1}\to\ell^{\infty}(\mathcal{S})$ from the proof of Proposition 4.6 does not.
+If the mapping $fs\mapsto M_{f}\lambda_{s}$ extended to an isomorphism between $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, then by Proposition 4.6, we would have $\mathbb{E}^{\prime}=\mathbb{E}$, a contradiction.
+∎
+We now arrive at the main theorem of this section.
+Theorem 4.8.
+Suppose $\mathcal{S}$ is a countable inverse semigroup with the property (FL) of Lledó and Martínez.
+If $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable (in the sense of [8]) if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
+Proof.
+By [14]*Theorem 4.18, $\textup{C}^{*}_{r}(\mathcal{S})$ is exact if and only if $\mathcal{R}_{\mathcal{S}}\cong\ell^{\infty}(\mathcal{S})\rtimes_{rr}\mathcal{S}$ is nuclear.
+By Theorem 4.7, this is equivalent to the nuclearity of $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}\cong\textup{C}^{*}_{r}(\mathcal{S}\ltimes\beta\mathcal{S})$.
+By [1]* Corollary 6.2.14, Theorem 3.3.7, this is equivalent to the amenability of $\mathcal{S}\ltimes\beta\mathcal{S}$.
+∎
+We conclude with a brief discussion.
+In light of the above theorem, the following question seems natural: What is the relationship between the amenability of $\mathcal{S}\ltimes\beta\mathcal{S}$ and the exactness of $\textup{C}^{*}_{r}(\mathcal{S})$ when $\mathcal{S}\ltimes\beta\mathcal{S}$ is non-Hausdorff?
+We remark that it was shown in [6]*Theorem A and independently in [3]*Theorem F that, for a (possibly non-Hausdorff) $\sigma$-compact étale groupoid $\mathcal{G}$, the amenability of $\mathcal{G}$ is equivalent to the nuclearity of $\textup{C}^{*}_{r}(\mathcal{G})$.
+Since $\mathcal{S}\ltimes\beta\mathcal{S}$ is $\sigma$-compact whenever $\mathcal{S}$ is countable, these results show that the action $\mathcal{S}\curvearrowright\beta\mathcal{S}$ is amenable in the sense of [8] if and only if the Buss-Exel-Meyer reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$ is nuclear.